# Loki: Low-Rank Keys for Efficient Sparse Attention

### v1

@@ Appendix E PCAAttn: Using reduced-dimensional attention scores directly in transformer models
| Llama2-7B | Exact TopK | 50% | - | 5.1191 | 57.2 | 68.9 | 28.3 | 31.2 | 63.9 | 86.0 |
| H2O | 50% | - | 5.1456 | 55.5 | 61.8 | 24.4 | 27.4 | 62.8 | 77.0 | |
| PCAAttn | - | 50% | 38.3997 | 33.3 | 53.2 | 21.7 | 14.2 | 50.5 | 73 | |
| Llama2-7B | Exact TopK | 25% | - | 5.1799 | 56.9 | 68.6 | 29.4 | 29 | 66.4 | 76.0 |
| H2O | 25% | - | 5.2809 | 50.1 | 51.6 | 21.1 | 17.8 | 55.2 | 55.0 | |
| PCAAttn | - | 25% | 243.2631 | 26.9 | 48.5 | 20.5 | 11.4 | 49.1 | 65.0 | |
@@
| Mistral-7B | Exact TopK | 50% | - | 4.9143 | 61.1 | 73.8 | 35.6 | 32.6 | 65.3 | 92.0 |
| H2O | 50% | - | 4.9560 | 59.4 | 58.6 | 26.4 | 23.0 | 62.4 | 71.0 | |
| PCAAttn | - | 50% | 396.8967 | 31.4 | 50.4 | 22.5 | 15.6 | 53.4 | 72.0 | |
| Mistral-7B | Exact TopK | 25% | - | 4.9170 | 60.4 | 73.0 | 35.4 | 30.0 | 65.3 | 85.0 |
| H2O | - | 5.0805 | 52.7 | 49.7 | 21.9 | 17.4 | 52.0 | 56.0 | ||
| PCAAttn | - | 25% | 933.6016 | 27.2 | 52.2 | 21.6 | 13.6 | 53.0 | 63.0 |
Table 6 shows the performance of PCAAttn on Llama2-7B and Mistral-7B models. We can see that our PCAAttn method performs poorly compared to all the baselines and the H2O method for all cache configurations. We believe that this happens because the application of rotary embeddings increases the dimensionality of the key vectors and using reduced dimensionality to store the keys results in loss of information. To further investigate this, let us look at Figure 9 which shows the rank at 90% explained variance for the key vectors across all layers and heads. Even though, the average rank per layer is around 50% of the full dimensionality, the rank for some layers and especially some heads within each layer is much higher. Due to the poor performance of PCAAttn, we do not include it in the final results and decide to focus on Loki instead in the main paper.

### v2

@@ Appendix E PCAAttn: Directly Leveraging Reduced-Dimensional Attention Scores
| Llama2-7B | Exact TopK | 0.5 | - | 5.1191 | 57.2 | 68.9 | 28.3 | 31.2 | 63.9 | 86.0 |
| H2O | 0.5 | - | 5.1456 | 55.5 | 61.8 | 24.4 | 27.4 | 62.8 | 77.0 | |
| PCAAttn | - | 0.5 | 38.3997 | 33.3 | 53.2 | 21.7 | 14.2 | 50.5 | 73 | |
| Llama2-7B | Exact TopK | 0.25 | - | 5.1799 | 56.9 | 68.6 | 29.4 | 29 | 66.4 | 76.0 |
| H2O | 0.25 | - | 5.2809 | 50.1 | 51.6 | 21.1 | 17.8 | 55.2 | 55.0 | |
| PCAAttn | - | 0.25 | 243.2631 | 26.9 | 48.5 | 20.5 | 11.4 | 49.1 | 65.0 | |
@@
| Mistral-7B | Exact TopK | 0.5 | - | 4.9143 | 61.1 | 73.8 | 35.6 | 32.6 | 65.3 | 92.0 |
| H2O | 0.5 | - | 4.9560 | 59.4 | 58.6 | 26.4 | 23.0 | 62.4 | 71.0 | |
| PCAAttn | - | 0.5 | 396.8967 | 31.4 | 50.4 | 22.5 | 15.6 | 53.4 | 72.0 | |
| Mistral-7B | Exact TopK | 0.25 | - | 4.9170 | 60.4 | 73.0 | 35.4 | 30.0 | 65.3 | 85.0 |
| H2O | 0.25 | - | 5.0805 | 52.7 | 49.7 | 21.9 | 17.4 | 52.0 | 56.0 | |
| PCAAttn | - | 0.25 | 933.6016 | 27.2 | 52.2 | 21.6 | 13.6 | 53.0 | 63.0 |
Table 5 shows the performance of PCAAttn on Llama2-7B and Mistral-7B models. We can see that our PCAAttn method performs poorly compared to all the baselines and the H2O method for all cache configurations. We believe that this happens because the application of rotary embeddings increases the dimensionality of the key vectors and using reduced dimensionality to store the keys results in loss of information. To further investigate this, we look back at Figure 10 which shows the rank at 90% explained variance for the key vectors across all layers and heads. Even though, the average rank per layer is around 50% of the full dimensionality, the rank for some layers and especially some heads within each layer is much higher. Due to the poor performance of PCAAttn, we do not include it in the final results and decide to focus on Loki instead in the main paper. Note, that we only tried the post-rotary transformations in PCAAttn, and it is possible that pre-rotary transformations might perform better, but we leave this for future work.